The incubation time for Rhode Island Red chicks is normally distributed with a mean of 21 days (μ) and standard deviation (σ) of approximately 1 day. If 1000 eggs are being incubated, how many chicks do we expect will hatch?
(a) In 19 to 23 days?
(b) 21 days or few?
Format that I can understand is below:
Z = x–μ =
σ
Change (Z to X)(To checking and verifying information)
X = Z σ + μ X = (?.00)(1) + 21 (X =??)


Answer by stanbon(75887)

The incubation time for Rhode Island Red chicks is normally distributed with a mean of 21 days (μ) and standard deviation (σ) of approximately 1 day. If 1000 eggs are being incubated, how many chicks do we expect will hatch?
(a) In 19 to 23 days?
z(19) = (19-21)/1 = -2
z(23) = (23-21)/1 = 2
P(19 <= x <= 23) = P(-2 <= z <= 2) = 0.9545
--------------------------------
(b) 21 days or few?
P(x <= 21) = P(z <= 2) = normalcdf(-100,2) = 0.9773
====================
Cheers,
Stan H.
=========================


Format that I can understand is below:
Z = x–μ =
σ
Change (Z to X)(To checking and verifying information)
X = Z σ + μ X = (?.00)(1) + 21 (X =??)